OpenStudy (anonymous):

James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done

OpenStudy (anonymous):

@Hero please come and help me on this, I really don't understand :(

hero (hero):

I don't see James' work?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

James added 3 and two-thirds plus 4 and one-eighth. His work is shown here. What did James do right? What did he do wrong? What should he have done? three and two-thirds plus 4 and one-eighth equals 7 and three-elevenths

OpenStudy (anonymous):

@Hero where did u go?

OpenStudy (anonymous):

that is the "work"

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

hi

hero (hero):

@carolinagb02, what do you think James did wrong?

OpenStudy (anonymous):

hey you asked me that i gave you the answer so did @Chineseboy15

OpenStudy (anonymous):

no u did not!, let me work with @hero please,

OpenStudy (anonymous):

he did not add correctly

OpenStudy (anonymous):

the answer should have been 7 19/24

OpenStudy (anonymous):

that's what i said

OpenStudy (anonymous):

I know but I want to know how to say the whole process in words, like in a paragraph!

OpenStudy (anonymous):

(example)

OpenStudy (anonymous):

I need to know what part of the adding he did wrong so I can say it in words!

OpenStudy (texaschic101):

3 2/3 + 4 1/8 dealing with just the fractions I get 2/3 + 1/8 = 16/24 + 3/24 = 19/24. dealing with the whole numbers I get 3 + 4 = 7 put it all together and you get : 7 19/24

OpenStudy (anonymous):

YOU SEE! that is the right answer thank you @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (anonymous):



OpenStudy (anonymous):

@carolinagb02 you can still ask me questions i'm sorry

OpenStudy (anonymous):



OpenStudy (texaschic101):

If I had $1 for every mistake I made, I would be rich :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

LOL!

OpenStudy (anonymous):

ok guys.

OpenStudy (anonymous):

thx for ur help, see u later!